Solve the following equations. 
x – 2 = 7
Solve the following equations. 
x – 2 = 7

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

x–2=7
x=7+2
x=9

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
X - 2 = 7
X = 7 + 2
X = 9
 
Hope it helps!
Thank you ✌️